QUESTION 6 The reaction 21Br(g) 리2(g) + Br2(g) Has an equilibrium constant K 0.012 If the initial concentration of IBr is 0.026, what is the equilibrium concentration of 12(g)? Enter your answer with 2 significant figures. QUESTION 7 For the reaction Initially the concentration of A was 0.2M. After reaching equilibrium, it was found that the concentration of A had decreased by 0.09. What is the value of the equilibrium constant K for this reaction? Enter your answer with 2 significant figures.
Answer #1

Question6

Answer

[I2] = 0.0023

Explanation

2IBr(g) < - - - - - > I2(g) + Br2(g)

K = [I2] [Br2] / [IBr]2 = 0.012

at equilibrium

[I2] = x

[Br2] = x

[IBr] = 0.026 - 2x

Therefore,

x2/(0.026 - 2x)2 = 0.012

x/(0.026 - 2x) = 0.10954

x + 0. 21908x = 0.002848

1.21908x = 0.002848

x = 0.002336

[ I2] = 0.0023

Question7

2A <- - - - - - > 3B

K = [A]2/[B]3

   at equlibrium

   [A] = 0.2 - 2x

   [B] = 3x

At equillibrium [A] = 0.2 - 0.09 = 0.11

   0.2 - 2x = 0.11

   - 0.2x = - 0.09

   x = 0.45

[B] = 3x = 3*0.45 = 1.35

K = (1.35)3/(0.11)2

= 203